# Invariant Coordinate Selection and Fisher discriminant subspace beyond the case of two groups

### v1

@@ Abstract
Invariant Coordinate Selection (ICS) is a multivariate technique that relies on the simultaneous diagonalization of two scatter matrices. It serves various purposes, including its use as a dimension reduction tool prior to clustering or outlier detection. Unlike methods such as Principal Component Analysis, ICS has a theoretical foundation that explains why and when the identified subspace should contain relevant information. These general results have been examined in detail primarily for specific scatter combinations within a two-cluster framework. In this study, we expand these investigations to include more clusters and scatter combinations. The case of three clusters in particular is studied at length. Based on these expanded theoretical insights and supported by numerical studies, we conclude that ICS is indeed suitable for recovering Fisher’s discriminant subspace under very general settings and cases of failure seem rare.
Keywords Dimension reduction  $\cdot$ Simultaneous diagonalization  $\cdot$ Mixture of elliptical distributions  $\cdot$ Scatter matrix  $\cdot$ Subspace estimation
@@ 1 Introduction
Recent research has sought to provide justifications for using these methods as preprocessing steps, often considering specific settings. The most popular framework involves the Gaussian mixture model, with the benchmark being whether the DR method can estimate Fisher’s discriminant subspace (FDS) Fisher (1936) in an unsupervised, or blind, manner. Outlier detection approaches are incorporated into this framework when some clusters are rare.
In the context of PCA, also known as tandem PCA Arabie and Hubert (1994) when it is used in conjunction with clustering, it is well-documented that FDS is seldom obtained Radojičić et al. (2021). PP is considered here when skewness or kurtosis are used as PP indices, and conditions under which these estimate the FDS are detailed in Radojičić et al. (2021) in the two-cluster case. ICS is the most recent DR method among those mentioned. ICS simultaneously diagonalizes two scatter matrices and has been considered for outlier detection in Archimbaud et al. (2018) and as tandem clustering with ICS in Alfons et al. (2024). Tyler et al. (2009) actually provide very general results for ICS with respect to FDS estimation. However, detailed results exist only for very specific scatter combinations in very specific settings. The goal of this paper is to extend these results to broader settings. The structure of the paper is as follows. In Section 2 we recall ICS in more detail, particularly focusing on its application in estimating FDS. Section 3 examines the properties of ICS for DR when the clusters in the mixtures exhibit no variability. In this context, and when the dimension of the FDS is equal to the number of groups minus one, we prove that for any scatter combination, the ICS eigenvalues do not depend on the cluster locations but only on the cluster proportions. Section 4 then gives closer attention to this model for a specific scatter combination and studies the behavior of the ICS eigenvalues when the cluster proportions vary. Section 5 explores the same scatter combination within a certain Gaussian mixture model consisting of three clusters. Section 6 validates the theoretical results and investigates the behavior of previously unexamined scatter combinations in such settings through a simulation study. The paper concludes with Section 7. Calculations and proofs are provided in the appendix together with some details on the parameters for our experiments.
@@ 2.1 General principle of ICS
Following the notations in Tyler et al. (2009), let $Y$ be a random vector of dimension $p$, with distribution function $F_{Y}$. Let $\mathcal{P}_{p}$ be the set of all positive definite symmetric matrices of order $p$. A scatter matrix of $Y$, denoted by $V(F_{Y})$, is a function of the distribution of $Y$ that is uniquely defined at $F_{Y}$, is such that $V(F_{Y})$ belongs to $\mathcal{P}_{p}$, and is affine equivariant, i.e. $V(F_{AY+b})=AV(F_{Y})A^{\top}$ for all non-singular matrices $A$ of dimension $p\times p$ and for all $b\in\mathbb{R}^{p}$, and where denotes the transpose operation. In what follows, we will drop the dependence on $F_{Y}$ and simply denote by $V$ the scatter $V(F_{Y})$ when the context is obvious.
@@
If the data distribution is elliptical, then all affine equivariant scatter matrices are proportional Nordhausen and Tyler (2015). This property is not true in general, outside the context of elliptical distributions, and ICS exploits the difference between two scatter matrices to detect non-elliptical structures such as clusters (Alfons et al., 2024) or outliers (Archimbaud et al., 2018). To perform this comparison, ICS relies on the simultaneous diagonalization of two scatter matrices $\mathrm{V}_{1}$ and $\mathrm{V}_{2}$ :
$H^{\top}\mathrm{V}_{1}H=D_{1}\quad\mbox{and}\quad H^{\top}\mathrm{V}_{2}H=D_{2},$
@@
where $D_{1}$ and $D_{2}$ are diagonal matrices such that $D_{1}^{-1}D_{2}=\mathrm{diag}(\rho_{1},\ldots,\rho_{p})$, $\rho_{1}\geq\ldots\geq\rho_{p}$ being the eigenvalues of $\mathrm{V}_{1}^{-1}\mathrm{V}_{2}$ sorted in descending order, and $H=(h_{1},\ldots,h_{p})$ is a $p\times p$ non singular matrix containing the corresponding eigenvectors. Usually $H$ is scaled such that $D_{1}=I_{p}$, where $I_{p}$ denotes the identity matrix of dimension $p$. The term “generalized eigendecomposition" (and correspondingly, “generalized eigenvalue" and “generalized eigenvector") is sometimes used in this context, but for simplicity, we avoid this terminology in the present paper.
@@
The $Z=H^{\top}Y$ transformation of $Y$ leads to new variables, that are invariant under affine transformation in the sense of Theorems 1 and 2 in Tyler et al. (2009), and are called invariant coordinates or components. Apart from being invariant by affine transformations, ICS has many applications and useful properties, as investigated for example in Nordhausen et al. (2008); Peña et al. (2010); Nordhausen et al. (2011); Loperfido (2013); Alashwali and Kent (2016); Loperfido (2021); Nordhausen and Ruiz-Gazen (2022); Archimbaud et al. (2023a). Of particular interest to us, is the following property for finite mixture models of the form:
$f_{Y}(y)=\mbox{det}(\Gamma)^{-1/2}\sum_{j=1}^{k}\alpha_{j}g_{j}((y-\mu_{j})^{\top}\Gamma^{-1}(y-\mu_{j})),$ (1)
@@
where $\alpha_{j}>0$ for $j\in\{1,\ldots,k\}$, with $\sum_{j=1}^{k}\alpha_{j}=1$, are the mixture proportions, the $\mu_{j}$ are distinct location vectors (also called group centers or group means), $\Gamma\in\mathcal{P}_{p}$ is the within-group scatter matrix parameter, and $g_{1},\ldots,g_{k}$ are non-negative functions. Thus, each of the $k$ components of the mixture is elliptical and the standard Gaussian mixture model is a special case. For such a model, Theorem 4 from Tyler et al. (2009) is fundamental and justifies the use of ICS as a dimension reduction method when the objective is clustering or anomaly detection. In short, this result proves that for model (1) and under some conditions, the ICS components associated with the largest and/or smallest eigenvalues span the FDS which is the one obtained by using the linear discriminant analysis in a supervized context. In the next subsection, we recall this theorem and also discuss some of its limitations.
@@ 2.2 ICS and the Fisher discriminant subspace
In a supervized context, when the distribution of $Y$ is a mixture of distributions, Fisher (1936) suggested to look for the linear function of the $p$ variables in $Y$ which maximize the ratio of the between-group variability to the within-group variability. This function corresponds to the projection of $Y$ onto the eigenvector associated with the largest eigenvalue of the simultaneous decomposition of the between and within-group covariance matrices. In the case of a mixture of two Gaussian groups with means $\mu_{1}$ and $\mu_{2}$ and equal within-group covariance matrices $\Gamma$, this eigenvector is equal to $\Gamma^{-1}(\mu_{1}-\mu_{2})$ and spans the FDS with dimension $q=k-1=1$. For a mixture of $k$ groups such as model (1), the FDS is spanned by the vectors $\Gamma^{-1}(\mu_{j}-\mu_{k})$ for $j\in\{1,\ldots,k-1\}$.
@@
Let us denote by $q$ the dimension of the vector space associated with the affine space spanned by the group centers $\mu_{j}$, $j\in\{1,\ldots,k\}$. According to Theorem 4 from Tyler et al. (2009), the simultaneous diagonalization of two scatter matrices results in at least one eigenvalue, denoted by $\rho*$, with a multiplicity greater than or equal to $p-q$. This eigenvalue is associated with an eigenspace of dimension at least $p-q$ which is the direct sum of the complementary space of the FDS (with dimension $p-q$) and, if the multiplicity of $\rho*$ is strictly larger than $p-q$, a space that is included in the FDS. If no eigenvalue has multiplicity greater than $p-q$, the subspace spanned by the eigenvectors associated with the other eigenvalues than $\rho*$ (larger or smaller) is the FDS.
@@
Theorem 4 in Tyler et al. (2009) is proven by using the equivariance property of the scatter matrices $\mathrm{V}_{1}$ and $\mathrm{V}_{2}$, and transforming the random vector $Y$, with distribution given by (1), in the following way. Let $M=(\mu_{1},\ldots,\mu_{k})$ be the matrix of rank $q$ which contains the location vectors of $Y$, and let $M_{0}=\Gamma^{-\frac{1}{2}}(M-\mu_{k}1_{k}^{\top})$, where $1_{k}$ is a $k$-dimensional vector of ones. The QR decomposition of $M_{0}$ is:
$M_{0}=PT=P\begin{pmatrix}T_{u}&0\\ 0&0\\ \end{pmatrix},$
@@
where $P$ is an orthogonal matrix, $T=(t_{1},\ldots,t_{k})$ with $t_{j}$ a $p$-dimensional vector for $j\in\{1,\ldots,k\}$, and $T_{u}$ is an upper triangular matrix of dimension $k-1\geq 1$ such that the last $k-1-q\geq 0$ rows are zero. Note that the dimension of $T_{u}$ in Tyler et al. (2009) differs slightly from ours, but this does not impact the remainder of the proof. The distribution of the transformed random vector $X=P^{\top}\Gamma^{-\frac{1}{2}}(Y-\mu_{k}\mathbf{1}_{k}^{\top})$ is a mixture of spherical distributions with density:
$f_{X}(x)=\sum_{j=1}^{k}\alpha_{j}g_{j}((x-t_{j})^{\top}(x-t_{j})),$ (2)
@@
where for $j\in\{1,\ldots,k\}$, $\alpha_{j}>0$, $\sum_{j=1}^{k}\alpha_{j}=1$, the $t_{j}$ are distinct, $t_{k}$ is the zero vector, and $g_{1},\ldots,g_{k}$ are non-negative functions. By decomposing the scatter matrices $\mathrm{V}_{1}(F_{X})$ and $\mathrm{V}_{2}(F_{X})$ in four blocks where the left top block has dimension $q\times q$, and using the affine equivariance of the scatter matrices, Tyler et al. (2009) are able to write:
$\mathrm{V}_{1}(F_{X})^{-1}\mathrm{V}_{2}(F_{X})=\left(\begin{array}[]{cc}A_{q}&0\\ 0&\gamma I_{p-q}\end{array}\right),$
@@
for some $\gamma>0$. This result implies that $\mathrm{V}_{1}(F_{X})^{-1}\mathrm{V}_{2}(F_{X})$ has an eigenvalue equal to $\gamma$ with multiplicity at least equal to $p-q$. The associated eigenspace can be written as the direct sum of the complementary space of the FDS (with dimension $p-q$) and, if the multiplicity of $\gamma$ is strictly larger than $p-q$, a space that is included in the FDS. The result is then derived for $Y$ by noting that the eigenvalues of $\mathrm{V}_{1}(F_{Y})^{-1}\mathrm{V}_{2}(F_{Y})$ are proportional to the ones of $\mathrm{V}_{1}(F_{X})^{-1}\mathrm{V}_{2}(F_{X})$. Thus there exists an eigenvalue $\rho*$ of $\mathrm{V}_{1}(F_{Y})^{-1}\mathrm{V}_{2}(F_{Y})$ with multiplicity at least $p-q$. Moreover, if the multiplicity of $\rho*$ is $p-q$, it is proven that the eigenvectors of $\mathrm{V}_{1}(F_{Y})^{-1}\mathrm{V}_{2}(F_{Y})$ which are not associated with $\rho*$ span the same subspace that is spanned by $\Gamma^{-1}(M-\mu_{k}1_{k}^{\top})$ and which corresponds to the FDS.
@@
This result holds for model (1) and for any pair of scatter matrices, making it very general. The problem, however, is that in order to perform dimension reduction with ICS, the eigenvalues associated with the FDS should be distinct from $\rho*$ meaning that the multiplicity of $\rho*$ should not be greater than $p-q$. Tyler et al. (2009) mention that this condition generally holds except for special cases, and that these special cases depend on the scatter pair and on the model parameters. In the following subsection, we recall some special cases for which the behavior of ICS is understood more precisely.

### v2

@@ Abstract
Invariant Coordinate Selection (ICS) is a multivariate technique that relies on the simultaneous diagonalization of two scatter matrices. It serves various purposes, including its use as a dimension reduction tool prior to clustering or outlier detection. ICS’s theoretical foundation establishes why and when the identified subspace should contain relevant information by demonstrating its connection with the Fisher discriminant subspace (FDS). These general results have been examined in detail primarily for specific scatter combinations within a two-cluster framework. In this study, we expand these investigations to include more clusters and scatter combinations. Our analysis reveals the importance of distinguishing whether the group centers matrix has full rank. In the full-rank case, we establish deeper connections between ICS and FDS. We provide a detailed study of these relationships for three clusters when the group centers matrix has full rank and when it does not. Based on these expanded theoretical insights and supported by numerical studies, we conclude that ICS is indeed suitable for recovering the FDS under very general settings and cases of failure seem rare.
Keywords Dimension reduction  $\cdot$ Mixture of elliptical distributions  $\cdot$ Scatter matrix  $\cdot$ Simultaneous diagonalization  $\cdot$ Subspace estimation
@@ 1 Introduction
Recent research has sought to provide justifications for using these methods as preprocessing steps, often considering specific settings. The most popular framework involves the Gaussian mixture model, with the benchmark being whether the DR method can estimate Fisher discriminant subspace (FDS) Fisher (1936) in an unsupervized, or blind, manner. Outlier detection approaches are incorporated into this framework when some clusters are rare.
The use of PCA in the context of clustering, also known as tandem PCA Arabie and Hubert (1994), can be justified using its relationship to principle points as pointed out in Tarpey et al. (1995); Matsuura and Kurata (2011). PCA will, however, only rarely recover the FDS as shown in Radojičić et al. (2021). PP is considered in this framework when skewness or kurtosis are used as PP indices, and conditions under which these estimate the FDS are detailed in Radojičić et al. (2021, 2025) in the two-cluster case. ICS is the most recent DR method among those mentioned. ICS simultaneously diagonalizes two scatter matrices and has been considered for outlier detection in Archimbaud et al. (2018) and for tandem clustering in Alfons et al. (2024). In Tyler et al. (2009), the authors actually provide very general results for ICS with respect to FDS estimation. However, detailed results exist only for very specific scatter combinations in very specific settings. The goal of this paper is to extend these results to broader settings. The structure of the paper is as follows. In Section 2 we recall ICS in more detail, particularly focusing on its application in estimating FDS. Section 3 examines the properties of ICS for DR when the clusters in the mixtures exhibit no variability. In this context, and when the dimension of the FDS is equal to the number of groups minus one, we prove that for any scatter combination, the ICS eigenvalues do not depend on the cluster locations but only on the cluster proportions. Section 4 then gives closer attention to this model for a specific scatter combination and studies the behavior of the ICS eigenvalues when the cluster proportions vary. Section 5 explores the same scatter combination within a Gaussian mixture model consisting of three clusters with a group centers matrix of rank one while the maximum rank is two. Through a simulation study and empirical application (Sections 6 and 7), we validate our theoretical results and examine the behavior of previously unstudied scatter combinations in settings studied in the previous sections. The paper concludes with Section 8. Calculations and proofs are provided in the appendix while some details on the parameters for our experiments are provided in the supplementary material.
@@ 2.1 General principle of ICS
Following the notations in Tyler et al. (2009), let $\mathbf{Y}$ be a random vector of dimension $p$, with distribution function $F_{\mathbf{Y}}$. Let $\mathcal{P}_{p}$ be the set of all positive definite symmetric matrices of order $p$. A scatter matrix of $\mathbf{Y}$, denoted by $\mathbf{V}(F_{\mathbf{Y}})$, is a function of the distribution of $\mathbf{Y}$ that is uniquely defined at $F_{\mathbf{Y}}$, is such that $\mathbf{V}(F_{\mathbf{Y}})$ belongs to $\mathcal{P}_{p}$, and is affine equivariant, i.e., $\mathbf{V}(F_{\mathbf{AY}+b})=\mathbf{AV}(F_{\mathbf{Y}})\mathbf{A}^{\top}$ for all non-singular matrices $\mathbf{A}$ of dimension $p\times p$ and for all $b\in\mathbb{R}^{p}$, and where denotes the transpose operation. In what follows, we will drop the dependence on $F_{\mathbf{Y}}$ and simply denote by $\mathbf{V}$ the scatter $\mathbf{V}(F_{\mathbf{Y}})$ when the context is obvious.
@@
If the data distribution is elliptical, then all affine equivariant scatter matrices are proportional Nordhausen and Tyler (2015). This property is not true in general, outside the context of elliptical distributions, and ICS exploits the difference between two scatter matrices to detect non-elliptical structures such as clusters (Alfons et al., 2024) or outliers (Archimbaud et al., 2018). To perform this comparison, ICS relies on the simultaneous diagonalization of two scatter matrices $\mathbf{V}_{1}$ and $\mathbf{V}_{2}$ :
$\mathbf{H}^{\top}\mathbf{V}_{1}\mathbf{H}=\mathbf{D}_{1}\quad\mbox{and}\quad\mathbf{H}^{\top}\mathbf{V}_{2}\mathbf{H}=\mathbf{D}_{2},$
@@
where $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$ are diagonal matrices such that $\mathbf{D}_{1}^{-1}\mathbf{D}_{2}=\mathrm{diag}(\rho_{1},\ldots,\rho_{p})$, $\rho_{1}\geq\cdots\geq\rho_{p}$ being the eigenvalues of $\mathbf{V}_{1}^{-1}\mathbf{V}_{2}$ sorted in descending order, and $\mathbf{H}=(\mathbf{h}_{1},\ldots,\mathbf{h}_{p})$ is a $p\times p$ non singular matrix containing the corresponding eigenvectors. Usually $\mathbf{H}$ is scaled such that $\mathbf{D}_{1}=\mathbf{I}_{p}$, where $\mathbf{I}_{p}$ denotes the identity matrix of dimension $p$. The term “generalized eigendecomposition" (and correspondingly, “generalized eigenvalue" and “generalized eigenvector") is sometimes used in this context, but for simplicity, we avoid this terminology in the present paper.
@@
The $\mathbf{Z}=\mathbf{H}^{\top}\mathbf{Y}$ transformation of $\mathbf{Y}$ leads to new variables, that are invariant under affine transformation in the sense of Theorems 1 and 2 in Tyler et al. (2009), and are called invariant coordinates or components. Apart from being invariant by affine transformations, ICS has many applications and useful properties, as investigated for example in Nordhausen et al. (2008); Peña et al. (2010); Nordhausen et al. (2011); Loperfido (2013); Alashwali and Kent (2016); Loperfido (2021); Nordhausen and Ruiz-Gazen (2022); Archimbaud et al. (2023a). Of particular interest to us, is the following property for finite mixture models of the form:
$f_{\mathbf{Y}}(\mathbf{y})=\mbox{det}(\boldsymbol{\Gamma})^{-1/2}\sum_{j=1}^{k}\alpha_{j}g_{j}((\mathbf{y}-\boldsymbol{\mu}_{j})^{\top}\boldsymbol{\Gamma}^{-1}(\mathbf{y}-\boldsymbol{\mu}_{j})),$ (1)
@@
where $\alpha_{j}>0$ for $j\in\{1,\ldots,k\}$, with $\sum_{j=1}^{k}\alpha_{j}=1$, are the mixture proportions, the $\boldsymbol{\mu}_{j}$ are distinct location vectors (also called group centers or group means), $\boldsymbol{\Gamma}\in\mathcal{P}_{p}$ is the within-group scatter matrix parameter, and $g_{1},\ldots,g_{k}$ are non-negative functions. Thus, each of the $k$ components of the mixture is elliptical and the standard Gaussian mixture model is a special case. For such a model, Theorem 4 from Tyler et al. (2009) is fundamental and justifies the use of ICS as a dimension reduction method when the objective is clustering or anomaly detection. In short, this result proves that for model (1) and under some conditions, the ICS components associated with the largest and/or smallest eigenvalues span the FDS which is the one obtained by using the linear discriminant analysis in a supervized context. In the next subsection, we recall this theorem and also discuss some of its limitations.
@@ 2.2 ICS and the Fisher discriminant subspace
In a supervized context, when the distribution of $\mathbf{Y}$ is a mixture of distributions, Fisher (1936) suggested to look for the linear function of the $p$ variables in $\mathbf{Y}$ which maximizes the ratio of the between-group variability to the within-group variability. This function corresponds to the projection of $\mathbf{Y}$ onto the eigenvector associated with the largest eigenvalue of the simultaneous decomposition of the between and within-group covariance matrices. In the case of a mixture of two Gaussian groups with means $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$ and equal within-group covariance matrices $\boldsymbol{\Gamma}$, this eigenvector is equal to $\boldsymbol{\Gamma}^{-1}(\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2})$ and spans the FDS with dimension $q=k-1=1$. For a mixture of $k$ groups such as model (1), the FDS is spanned by the vectors $\boldsymbol{\Gamma}^{-1}(\boldsymbol{\mu}_{j}-\boldsymbol{\mu}_{k})$ for $j\in\{1,\ldots,k-1\}$.
@@
Let us denote by $q$ the dimension of the vector space associated with the affine space spanned by the group centers $\boldsymbol{\mu}_{j}$, $j\in\{1,\ldots,k\}$. According to Theorem 4 from Tyler et al. (2009), the simultaneous diagonalization of two scatter matrices results in at least one eigenvalue, denoted by $\rho*$, with a multiplicity greater than or equal to $p-q$. This eigenvalue is associated with an eigenspace of dimension at least $p-q$ which is the direct sum of the complementary space of the FDS (with dimension $p-q$) and, if the multiplicity of $\rho*$ is strictly larger than $p-q$, a space that is included in the FDS. If no eigenvalue has multiplicity greater than $p-q$, the subspace spanned by the eigenvectors associated with the other eigenvalues than $\rho*$ (larger or smaller) is the FDS.
@@
Theorem 4 in Tyler et al. (2009) is proven by using the equivariance property of the scatter matrices $\mathbf{V}_{1}$ and $\mathbf{V}_{2}$, and transforming the random vector $\mathbf{Y}$, with distribution given by (1), in the following way. Let $\mathbf{M}=(\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{k})$ be the matrix of rank $q$ which contains the location vectors of $\mathbf{Y}$, and let $\mathbf{M}_{0}=\boldsymbol{\Gamma}^{-\frac{1}{2}}(\mathbf{M}-\boldsymbol{\mu}_{k}\mathbf{1}_{k}^{\top})$, where $\mathbf{1}_{k}$ is a $k$-dimensional vector of ones. The QR decomposition of $\mathbf{M}_{0}$ is:
$\mathbf{M}_{0}=\mathbf{PT}=\mathbf{P}\begin{pmatrix}\mathbf{T}_{u}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\\ \end{pmatrix},$
@@
where $\mathbf{P}$ is an orthogonal matrix, $\mathbf{T}=(\mathbf{t}_{1},\ldots,\mathbf{t}_{k})$ with $\mathbf{t}_{j}$ a $p$-dimensional vector for $j\in\{1,\ldots,k\}$, and $\mathbf{T}_{u}$ is an upper triangular matrix of dimension $k-1\geq 1$ such that the last $k-1-q\geq 0$ rows are zero. Note that the dimension of $\mathbf{T}_{u}$ in Tyler et al. (2009) differs slightly from ours, but this does not impact the remainder of the proof. The distribution of the transformed random vector $\mathbf{X}=\mathbf{P}^{\top}\boldsymbol{\Gamma}^{-\frac{1}{2}}(\mathbf{Y}-\boldsymbol{\mu}_{k}\mathbf{1}_{k}^{\top})$ is a mixture of spherical distributions with density:
$f_{\mathbf{X}}(\mathbf{x})=\sum_{j=1}^{k}\alpha_{j}g_{j}((\mathbf{x}-\mathbf{t}_{j})^{\top}(\mathbf{x}-\mathbf{t}_{j})),$ (2)
@@
where for $j\in\{1,\ldots,k\}$, $\alpha_{j}>0$, $\sum_{j=1}^{k}\alpha_{j}=1$, the $\mathbf{t}_{j}$ are distinct, $\mathbf{t}_{k}$ is the zero vector, and $g_{1},\ldots,g_{k}$ are non-negative functions. By decomposing the scatter matrices $\mathbf{V}_{1}(F_{\mathbf{X}})$ and $\mathbf{V}_{2}(F_{\mathbf{X}})$ in four blocks where the left top block has dimension $q\times q$, and using the affine equivariance of the scatter matrices, Tyler et al. (2009) are able to write:
$\mathbf{V}_{1}(F_{\mathbf{X}})^{-1}\mathbf{V}_{2}(F_{\mathbf{X}})=\left(\begin{array}[]{cc}\mathbf{A}_{q}&\mathbf{0}\\ \mathbf{0}&\gamma\mathbf{I}_{p-q}\end{array}\right),$
@@
for some $\gamma>0$. This result implies that $\mathbf{V}_{1}(F_{\mathbf{X}})^{-1}\mathbf{V}_{2}(F_{\mathbf{X}})$ has an eigenvalue equal to $\gamma$ with multiplicity at least equal to $p-q$. The associated eigenspace can be written as the direct sum of the complementary space of the FDS (with dimension $p-q$) and, if the multiplicity of $\gamma$ is strictly larger than $p-q$, a space that is included in the FDS. The result is then derived for $\mathbf{Y}$ by noting that the eigenvalues of $\mathbf{V}_{1}(F_{\mathbf{Y}})^{-1}\mathbf{V}_{2}(F_{\mathbf{Y}})$ are proportional to the ones of $\mathbf{V}_{1}(F_{\mathbf{X}})^{-1}\mathbf{V}_{2}(F_{\mathbf{X}})$. Thus there exists an eigenvalue $\rho*$ of $\mathbf{V}_{1}(F_{\mathbf{Y}})^{-1}\mathbf{V}_{2}(F_{\mathbf{Y}})$ with multiplicity at least $p-q$. Moreover, if the multiplicity of $\rho*$ is $p-q$, it is proven that the eigenvectors of $\mathbf{V}_{1}(F_{\mathbf{Y}})^{-1}\mathbf{V}_{2}(F_{\mathbf{Y}})$ which are not associated with $\rho*$ span the same subspace that is spanned by $\boldsymbol{\Gamma}^{-1}(\mathbf{M}-\boldsymbol{\mu}_{k}\mathbf{1}_{k}^{\top})$ and which corresponds to the FDS.
@@
This result holds for model (1) and for any pair of scatter matrices, making it very general. The problem, however, is that in order to perform dimension reduction with ICS, the eigenvalues associated with the FDS should be distinct from $\rho*$ meaning that the multiplicity of $\rho*$ should not be greater than $p-q$. In Tyler et al. (2009), the authors mention that this condition generally holds except for special cases, and that these special cases depend on the scatter pair and on the model parameters. In the following subsection, we recall some special cases for which the behavior of ICS is understood more precisely.